Problem 1 (a) Put x = -1 and deduce that -Ε0ω n n |(-1)6 0 = k k 0 |(b) By adding (11) and (1 - 1), deduce that Σ ). 2-1 2
Answer #1

that . We know the expansion of 7 = 4 711 20 + ( T+ x + + O=(1-1 ... (-U) @ Putting x=-1 O:(1-1)=0-1-()-))+...+(-5° 0)+...+© ) . 29=(1+0)95 nCo+n(y + 2 + . -tran, — 0 OPC (-1)9 6. Co 39(1°+ng 17! (13tnout na 60() Oºs n6 -nca + ncg ....+(-09.. Add